OpenStudy (anonymous):

A chain saw requires 4 hours of assembly and a wood chipper 6 hours. A maximum of 48 hours of assembly time is available. The profit is $150 on a chain saw and $220 on a chipper. How many of each should be assembled for maximum profit?? This is what I have... I don't know the process/ssteps to come to the answer: Let c= chipper, let s= saw obj: Profit=150s+220c const: c>=0, s>=0 4s+6c<=48

OpenStudy (anonymous):

is this being taught in calculus pr algebra II, there are a few ways to solve it. with calculus it is very easy

OpenStudy (anonymous):

Yes, calculus

OpenStudy (anonymous):

ok. so this is a standard optimization problem. start by isolating s or c

OpenStudy (anonymous):

do you know what i mean by that?

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

Yes I get that... Isolating s or c

OpenStudy (anonymous):

Would I isolate by kind of acting like i would solve for s or c

OpenStudy (anonymous):

That's what you mean right

OpenStudy (anonymous):

pick either s or c to isolate from the equation 4s +6c = 48

OpenStudy (anonymous):

6c=48-4s

OpenStudy (anonymous):

now divide both sides by 6...

OpenStudy (anonymous):

ok I got it 2

OpenStudy (anonymous):

C=8-.66s

OpenStudy (anonymous):

how do i know the min and find the max profit?

OpenStudy (anonymous):

now plug that in to your profit equation. now u can have everything in terms of s

OpenStudy (anonymous):

now look for all times when the derivative of profit equation equals 0

OpenStudy (anonymous):

you shouldnt have told me you were taking calculus. i started thinking about it like calculus. there is very easy way to do this. we now that 48 is divisible by 6 and it is also divisible by 4. so it is possible that you spend all your time on chainsaws or all your time on woodchippers. think about how much money you make per hour for each. for woodchipper its 220/6 =36.6 for chansaw its 150/4 = 37.5

OpenStudy (anonymous):

you will make more per hour is spend all your time making chainsaws because 37.5 > 36.666

OpenStudy (anonymous):

OOh arite. i gotcha. ... i missed the class when the professor was teaching this.. i was studying for another ex am =/

OpenStudy (anonymous):

would i do the same process for this one as well: deluxe coffee is to be mixed with regular coffee to make at least 50 pounds of a blended coffee. the mixture must contain at least 10 pounds of deluxe coffee. deluxe coffee cost $6 per pound and regular coffee $5 per pound. How many pounds of each kind of coffee should be used to minimize cost?

OpenStudy (anonymous):

this one gets a little bit trickier because 6 doesnt go into 50 nicely

OpenStudy (anonymous):

nevermind/ scratch that. its irelevant

OpenStudy (anonymous):

lol ok.scratched

OpenStudy (anonymous):

you want to make this as cheap as possible. but you need at least 10 pounds of deluxe. so use 10 pounds of deluxe. 10 x 6 = $60 dollars. for the other 40 pounds use regular. 40 x 5 = $200. total cost 60 + 200 = $260

OpenStudy (anonymous):

it seems so easy when you know how to sort the info

OpenStudy (anonymous):

so .. let x= deluxe, y=regular \[x+y \ge 50 \] constraints:\[x \ge60, y \ge40\] ?

OpenStudy (anonymous):

i need to: 1. label all variables 2. show the objective function 3. list constraints 4. sketch the region and label all vertices defined by constraints 5. find the vertex that either min or maxed the fucntion

OpenStudy (anonymous):

the constraints are x > or equal to 0 and y > or equal to 0

OpenStudy (anonymous):

objective: 6x+5y=260

OpenStudy (anonymous):

i isolated x and got x=4.3-3.3y .. is that my obj or did i have it correct above

OpenStudy (anonymous):

i think whats mean by objective is to find the minimum cost

OpenStudy (anonymous):

it hve to show an equation to minimize

OpenStudy (anonymous):

so i guess that would be correct. i just think its strange to put the resultant minimum price in the objectice

OpenStudy (anonymous):

idk my professor is a bit nutty... i must go to class now, thank so much for your help!